OpenStudy (anonymous):

What is true about the sum of the two polynomials? 6s2t – 2st2 4s2t – 3st2

OpenStudy (anonymous):

L=LYM+LHiggs+Llep.+Lqrk.+LYuk. where LYM=−14TrFabFab LHiggs=−(Daϕ†)(Daϕ)+14λ(ϕ†ϕ−12v2)2 Llep.=iℓ†I/DℓI+ie¯†I/De¯I Lqrk.=iq†I/DqI+iu¯†I/Du¯+id¯†I/Dd¯I LYuk.=−yIJϕℓIe¯J−y′IJϕqId¯J −y′′IJϕ†qIu¯J+h.c.